#22 Suppose cars pass an intersection at a rate of 120 per hour. What is the probability that we need to wait between 25 and 30 seconds for the next car to pass? Please round your answer to 4 decimal places. Expert Answer
Answer #1

interarrival time between cars =60*60/120 =30 seconds

for exponential distribution:

F(x)=P(X<x)=1-e-x/β

from above :

P(25<X<30)=(1-exp(-30/30)-(1-exp(-25/30))= 0.0667